which of the points following the ineqaulity satifes the below
y<=-x+6
(2,0), (-3,9), (-4,12)

Answer by sdmmadam@yahoo.com(530)

y<=-x+6 ----(1)
Consider (2,0) in (1)
That is put x = 2 and y = 0 in
y<=-x+6 ----(1)
we get
0 < = -2+6
0 < = 4
And since 0 < 4 we can take the strict inequality
--------------------------------------------------
Consider (-3,9) in (1)
That is put x = -3 and y = 9
y<=-x+6 ----(1)
we get
9 < = -(-3)+6
9 < = 3+6
And since 9 = 9
here the equality holds
-------------------------
Consider (-4,12)in (1)
That is put x = -4 and y = 12
y<=-x+6 ----(1)
we get
12 < = -(-4)+6
12 < = 4+6
And since 12 not less than or equal to 10
here the given INEQUALITY DOES NOT HOLD